OpenStudy (anonymous):

find a if sqrta(sqrt6)=3sqrt2 find b if sqrt10(sqrt5+sqrtb)=5sqrt2+2sqrt5

OpenStudy (zepp):

\[\large \sqrt{\sqrt{6}}=3\sqrt{2}\] Like this? D:

OpenStudy (anonymous):

no

OpenStudy (zepp):

\[\large \sqrt{a}\sqrt{6}=3\sqrt{2}\]\[\large \sqrt{10}(\sqrt{5}+\sqrt{b})=5\sqrt2+2\sqrt5\] Like this?

OpenStudy (anonymous):

\[Find a if \sqrt{a}(\sqrt{6})=3\sqrt{2}\]

OpenStudy (anonymous):

yes like that

OpenStudy (zepp):

Alright \[\large \sqrt{a}\sqrt{6}=3\sqrt{2}=\sqrt{9}\sqrt{2}=\sqrt{2*9}=\sqrt{18}\] \(\large \sqrt{a}\sqrt{6}=\sqrt{18}\rightarrow \sqrt{a}=\frac{\sqrt{18}}{\sqrt{6}}\rightarrow a=\sqrt{\frac{18}{6}}^2=\sqrt{3}^2=3\)

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

how did u know a was 9

OpenStudy (anonymous):

nvm i get that

OpenStudy (anonymous):

i dont get the last part

OpenStudy (zepp):

Which part?

OpenStudy (zepp):

The second line?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the answer not sqrt18

OpenStudy (zepp):

That was still the first question!

OpenStudy (anonymous):

its ok nvm i understand, what about the next one

OpenStudy (anonymous):

do i do the f2nd one like the first

OpenStudy (anonymous):

b=2 correct??

OpenStudy (zepp):

\[\large \begin{align} \sqrt{10}(\sqrt{5}+\sqrt{b})&=5\sqrt2+2\sqrt5 \\ \sqrt{10*5}+\sqrt{10b}&=\sqrt{25}\sqrt{2}+\sqrt{4}\sqrt{5}\\ \sqrt{50}+\sqrt{10b}&=\sqrt{50}+\sqrt{20}\\ \sqrt{50}+\sqrt{10b}-\sqrt{50}&=\sqrt{20}\\ \sqrt{10}*\sqrt{b}&=\sqrt{20}\\ \sqrt{b}&=\sqrt{\frac{20}{10}}\\ \sqrt{b}&=\sqrt{2}\\ b&=2 \end{align}\]

OpenStudy (zepp):

Correct! :)

OpenStudy (anonymous):

thanks